OpenStudy (anonymous):

[4 - (3c - 1)][6 - ( 3c - 1)] another type of problem that i dont understand

OpenStudy (anonymous):

[4 - (3c - 1)][6 - ( 3c - 1)] (4-3c+1)(6-3c+1) 24-12c+4-18c+9c^2-3c+6-3c+1 9c^2-36c+31

OpenStudy (anonymous):

im studying for a test tommorow so that is why i keep asking for people to solve these im asking questions that i do not understand and thanks

OpenStudy (asnaseer):

the first step should be to simplify each bracketed term. so lets look at the first term: 4 - (3c - 1) = 4 - 3c + 1 = 5 - 3c do you understand this?

OpenStudy (asnaseer):

I can explain further if you don't understand, or I can move on to the next term if you do understand. I need to know first if you understood the first step?